Solve:
sqrt of 6x +8 +2x=16

Answer by mathchemprofessor(65)

sqrt of 6x +8 +2x=16
Squaring both sides, we get,
6x+8+2x = 16^2
8x+8 =256
8(x+1)=256
x+1=256/8
x+1=32
x=32-1=31
x=31